If Rachel invests $2700 today in an account that pays 6% compounded annually, how long will it take for her to accumulate $8,500 in her account?
Answer #1

Ans 19.68 years

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
8500 / 2700 = (1 + 6%)^n
3.1481481481 = 1.06^n
n = 19.68 years